Select the equation in which the graph of the line has a negative slope, and the y-intercept equals


5x + 3y = 4

5x - 3y = –4

4x + 16y = –20

6x + 2y = –8


Answer by Alan3354(69443)

Select the equation in which the graph of the line has a negative slope, and the y-intercept equals


5x + 3y = 4

5x - 3y = –4

4x + 16y = –20

6x + 2y = –8
======================
Solve for y
--> y = mx + b where m = slope